OpenStudy (anonymous):

Question 2.2. Which ordered pair makes the inequality true? 2x + y > 8 (Points : 1) (–13, 0) (–2, 13) (0, 4) (5, –3)

OpenStudy (vishweshshrimali5):

Best way would be evaluate Left hand side and right hand side of the inequality for each option.

OpenStudy (vishweshshrimali5):

I will give you an example: Take the first option -> (-13,0): LHS = 2x + y Now, here x = -13 and y = 0 So: 2x + y = 2(-13) + 0 = -26 RHS = 8 Now, is -26 > 8 ? NO!!! So, this pair does NOT make the inequality true.

OpenStudy (vishweshshrimali5):

Similarly, you can try for other options.

OpenStudy (vishweshshrimali5):

Can you tell me which would be the correct answer ?

OpenStudy (anonymous):

umm...(5,-3) ? Ah i suck at math lol

OpenStudy (anonymous):

@vishweshshrimali5 am I wrong?

OpenStudy (vishweshshrimali5):

Yeah :( See for (5,-3): 2y + x = (2 * 5) - 3 = 10 - 3 = 7 Now, is 7 > 8 ? NO !!

OpenStudy (vishweshshrimali5):

Try again :)

OpenStudy (anonymous):

(0,4)?

OpenStudy (vishweshshrimali5):

What is (2x+y) for (0,4) ?

OpenStudy (vishweshshrimali5):

x = 0 and y = 4

OpenStudy (anonymous):

x=-2y-1+-0.5y? that's what I got idk i have a hard time with stuff like this

OpenStudy (anonymous):

am I wrong again @vishweshshrimali5 ?

OpenStudy (vishweshshrimali5):

See: Lets not make it too difficult. We have: LHS = 2x + y y = 4 and x = 0 Now 2x + y = 2 * 0 + 4 = 4 Now 4 is NOT greater than 8 So, this is NOT the correct option

OpenStudy (anonymous):

so, then the only option left is (-2,13)

OpenStudy (vishweshshrimali5):

What is 2x + y for (-2,13)? x =-2 and y = 13 So: 2x + y = 2 * (-2) + 13 = -4 + 13 = 9 which is surely greater than 8 So, this is the correct answer